At the start of the season, two teams are thought to independently have a 40% chance and a 60% chance of making the final. What is the probability that neither team makes the final?
Answer by greenestamps(13334)



P(A) = 0.40, so P(not A) = 0.60

P(B) = 0.60, so P(not B) = 0.40

P((not A) AND (not B)) = P(not A) * P(not B) = 0.60*0.40 = 0.24